We denote by SF (x) the Fourier series corresponding to the even, periodic prolongation of period 4 of the function h (x). Without calculating the coefficients of SF (x), give the values of SF (−1.5) and SF (123). Justify your answers.
Answer #1

Even por preodic elongation: &(onl=hens h(u) = -25 rs- i -1 <n<1 relesa hlu+4)= h(n) KI 1<x< 2 SF (-1.5) - h(-1.5) = -(-1-5)